OpenStudy (anonymous):

write the expression in standard form 8-i / 8+i

OpenStudy (anonymous):

i of imaginary?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i got 7/2 - 9/2i as the answer

OpenStudy (anonymous):

is this right?

OpenStudy (tkhunny):

Remember your Order of Operations. It is ALWAYS working. You did try to get it with spacing, but it is much more clear with parentheses. (8-i)/(8+i)

OpenStudy (anonymous):

(8-i)/(8+i) is the question and my answer is 7/2 - 9/2i

OpenStudy (anonymous):

on the book it didn't have paranthesis

OpenStudy (tkhunny):

It very often doesn't IN A BOOK. You are not writing the same way, here. ALWAYS write what you mean and make sure it says what you want. Trust no one - book or otherwise. Strive for clarity.

OpenStudy (tkhunny):

Show numerator and denominator separately. \(\dfrac{8-i}{8+i}\cdot \dfrac{8-i}{8-i} = \)

OpenStudy (anonymous):

8-8i-i+i squared / 8-i+i-i squared

OpenStudy (anonymous):

then i got 8-9i+(-1) / 1-(-1)

OpenStudy (anonymous):

7-9i / 2 = 7/2-9/2i

OpenStudy (tkhunny):

No good. Please go over that again. \(\dfrac{8-i}{8+i} = \dfrac{8-i}{8+i}\cdot\dfrac{8-i}{8-i} = \dfrac{64 - 16i + i^{2}}{64 - i^{2}}\) Simplify and try again.